The radius of the outer sphere of a spherical capacitor is five times the radius of its 

inner shell. What are the dimensions of this capacitor if its capacitance is 5.00 pF?



1
Expert's answer
2022-03-09T10:33:08-0500

The capacitance of a spherical capacitor is given by

"C=\frac{1}{k\left(\frac{1}{R_1}-\frac{1}{R_2}\right)}"

Hence

"\frac{1}{R_1}-\frac{1}{R_2}=\frac{1}{kC}"

"\frac{1}{R_1}-\frac{1}{5R_1}=\frac{1}{kC}"

"\frac{4}{5R_1}=\frac{1}{kC}"

"R_1=\frac{4}{5}kC\\
=\frac{4}{5}*9*10^9*5.00*10^{-12}\\
=0.036\:\rm m=3.6\: cm"

"R_2=5R_1=5*3.6\rm\: cm=18\: cm"